OpenStudy (anonymous):

x^2 – 8x + 15 please explain how to factor this trinomial ?

OpenStudy (anonymous):

(x-?)(y-?)

OpenStudy (anonymous):

(x-3)(x-5)

OpenStudy (anonymous):

what 2 numbers when multiplied equal 15 the same numbers when added equal 8 (x-3)(x-5)

OpenStudy (anonymous):

-5*-3=15 -5+-3=-8

OpenStudy (anonymous):

Thank you :D

OpenStudy (anonymous):

a^2 – a – 20 what do you do now since theres no two numbers ..?

OpenStudy (anonymous):

There is....

OpenStudy (anonymous):

same thing. what 2 numbers when multiplied equal 20 the same numbers when added equal a which is actually just 1 5*-4=-20 4+-5=-1

OpenStudy (anonymous):

Oh yeah a is 1a

OpenStudy (anonymous):

Just for info, even though it is a trinomial, most people would call it a quadratic

OpenStudy (anonymous):

oops i mean 4*-5=20 (x-5)(x+4)

OpenStudy (anonymous):

a^2 – 5a – 20 Im sorry last one i cant figure it out ?